HUBBLE SPACE TELESCOPE DAILY REPORT #2466 PERIOD COVERED: 0000Z (UTC) 09/27/99 - 0000Z (UTC) 09/28/99 Daily Status Report as of 271/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8345 (Multipolar Bubbles and Jets in Very Low Excitation Planetary Nebulae - A WF/PC-2 H-Alpha Imaging Survey) The WF/PC-2 was used to make H-Alpha observations of the very low excitation planetary nebulai PK226-30D1 and PK007-04D1 in order to study their multipolar bubbles and jets. Comparison of the morphology of this young PNe sample with more evolved PNe populations will set new constraints on models for the production of jet-like outflows and nebular evolution. The observations were completed as planned, and no problems were reported. 1.2 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed Three Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.5 Completed Three Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 7279 (The Impact of Cluster Substructure on Galaxies: A Case Study in ABELL 2125) The WF/PC-2 was used to make observations of the Abell-2125 galactic cluster (a.k.a., the Butcher-Oemler Cluster) in order to study the impact of the substructure of the cluster on galaxies. The VLA data show that, as well as a large blue-galaxy fraction {0.19} for its redshift, it has an unprecedented number of radio galaxies {30 with confirmed redshifts}, evenly split between old stellar populations and starbursting galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 8412 (STIS CCD Imaging Flats C8) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field stability of the unit over a monthly period. A bias calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Three Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.9 Completed STIS/CCD 8411 (STIS CCD Spectroscopic Flats C-8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed STIS/CCD 8313 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of 9765 and its host galaxy. The propose is to obtain snapshot, unfiltered STIS images of distant galaxies of known redshift which have hosted supernovae {SNe} of Type IA found via the Supernova Cosmology Project. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed STIS/CCD 8271 (A Snapshot Survey of LMC Planetary Nebulae: A Study of Nebular and Late Stellar Evolution) The Space Telescope Imaging Spectrograph (CCD) was used to make medium-resolution slitless spectroscopy snapshot images of the planetary nebulae LMC-SMP-34 and LMC-SMP-46 located in the Large Magellanic Cloud to study the parallel evolution of the nebulae and their central stars. The observations were completed as planned, and no problems were reported. 1.12 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.13 Completed S/C 7725 (STIS NUV MAMA Anomalous Recovery) This proposal is designed to permit recovery of the NUV MAMA detector after an anomalous shutdown. The proposal was executed as scheduled, and no anomalies were noted. 1.14 Completed Three Sets of STIS/MA1 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 35 Successful: 35 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset twice. Per an operations request, the STIS MAMA-2 high voltage event flag was cleared at 270/1142Z as the first step in the STIS MAMA-2 recovery. Later, between 270/1635Z and 270/1740Z, the MAMA-2 low voltage was turned on and closely monitored via SSR dumps and analysis. More activities in this incremental process will occur today and on succeeding days. A TTR was written for a TDRS-W failure. Engineering data was lost from 270/221449Z until 270/225330Z. TDRS-W supports were changed to the SSA for the duration of the failure. The SSA transmitter was turned on at 270/1557Z and off at 271/03317Z per ROP per ROP SR-1. HST operations switched to the CCS “A” string using release 4.0, with CCS “G” string configured as the backup string. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of the incremental STIS MAMA-2 recovery process. Continuation of normal science observations and calibrations. /DMH